ordered pairs >
2y-5x=16
2x+3y=13

Answer by pradhyumna.kumar@gmail.com(62)

ordered pairs >
2y-5x=16
2x+3y=13 after solving the equations -5x+2y=16
2x+3y=13

we get x = -22/19
y = 97/19